# LATTE: Latent Trajectory Embedding for Diffusion-Generated Image Detection

### v1

@@ B.2 Influence of separate latent processing strategy
B.3 Effect of position encoding in CLS-pooling
Section 4 of the paper ablates the aggregation mode by considering a CLS-style pooling variant. Specifically, we leverage a self-attention mechanism over the sequence of latents, using a prepended learnable [CLS] token. The configuration evaluated in the main results includes learnable positional embeddings to provide the model with information about the order of latents along the trajectory. Here, we conduct a focused ablation to isolate the effect of the positional embeddings. Specifically, we compare the full model (“CLS-pooling w/ pos. enc.”) to a variant that uses the same CLS-based self-attention but omits positional embeddings (“CLS-pooling w/o pos. enc.”), removing any explicit indication of timestep order. As shown in Table 8, providing sequence order information results in a significant improvement of 7.2% in average accuracy, confirming that timestep position is an important signal when aggregating latents jointly.
Table 8: Accuracy(%) comparison between models with and without explicit sequence order. Explicit positional embeddings improve accuracy by 7.2% over the implicit variant, but still fall short of the baseline strategy.
Sequence order Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
CLS-pooling w/o pos. enc. 75.5 99.7 99.6 72.2 78.9 98.3 75.6 59.3 82.4
CLS-pooling w/ pos. enc. 85.8 99.0 99.0 83.2 86.7 96.8 88.7 77.8 89.6
@@ B.3 Effect of position encoding in CLS-pooling
Despite this gain, both CLS-based variants remain less effective than the default LATTE architecture, which processes each timestep independently and aggregates the outputs via average pooling. Interestingly, the "CLS-pooling w/ pos. enc." variant demonstrates better performance on certain subsets - 9.2% increase on ADM and 7.9% on VQDM - suggesting that this CLS-based design, paired with sequence order cues can be beneficial in specific contexts. However, on average, the baseline approach still outperforms both CLS-pooling variants, highlighting that preserving timestep-specific specialization is ultimately more effective.
B.4 Influence of vision backbone fine-tuning
Our default setup fine-tunes the vision encoder. To quantify the added benefit of this choice, we compare against a variant where we freeze the backbone and train only the LATTE-specific components. Table 9 reports per-generator accuracy for both settings.
Table 9: Accuracy (%) comparison for different vision backbones and pooling strategy combinations across generators in the GenImage dataset.
@@ B.4 Influence of vision backbone fine-tuning
We observe major improvements for both vision backbones when fine-tuned, with 15.5% accuracy gain for CLIP ViT-L/14 [48] and 9% for ConvNeXt [56] . This likely stems from the fact that frozen backbones retain features that were never explicitly optimized for real vs. fake discrimination, leading to an embedding space that is misaligned with the objectives of generated image detection. Without adaptation, our model may struggle to effectively ground latent trajectories in meaningful visual semantics. Fine-tuning, by contrast, enables the backbone to specialize its representations for this task, enhancing the alignment between visual and latent features essential for robust detection.
B.5 Influence of vision backbones across different processing strategies
In the main paper, we analyze the effect of two different vision backbones - ConvNeXt and CLIP ViT-L/14 - under our default setup (i.e. separate latent processing). In this section, we provide a more comprehensive analysis by exploring how backbone choice interacts with different latent processing strategies - namely, Separate (default), Joint (as introduced in Section A.2) and CLS-based (as described in Section 4.4) - to better understand their combined influence on performance.
Table 10: Detection accuracy (%) for different backbone and pooling strategy combinations across generators in the GenImage dataset.
Config. Backbone Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
Separate CLIP ViT-L/14 98.2 99.9 100 68.4 91.8 99.9 58.9 72.1 86.1
@@ B.5 Influence of vision backbones across different processing strategies
Joint CLIP ViT-L/14 96.2 100 100 63.4 85.2 99.9 59.2 65.9 83.7
@@
CLS CLIP ViT-L/14 93.3 99.9 99.9 64.7 85.9 99.4 58.3 63.0 83.0
[…]
As shown in Table 10, Separate latent processing remains the most effective strategy across both backbones, confirming the value of maintaining timestep-specific representations when modeling the denoising trajectory. Among the configurations, ConvNeXt [56] with Separate pooling achieves the highest overall accuracy (90.9%) and demonstrates stronger generalization to challenging generators like VQDM(80.8%) and BigGAN(89.7%). In contrast, CLIP ViT-L/14, while excelling on in-distribution generators (e.g., SDv1.4, SDv1.5, Wukong >99.9%), sees notable drops on out-of-distribution ones, suggesting a tendency to overfit to the training domain. Overall, the findings indicate that ConvNeXt generalizes more reliably across diverse generative styles, whereas CLIP ViT-L/14 is more prone to overfitting.
Appendix C Embedding Space Analysis
To complete our embedding space analysis from Section 4.5, Figure 8 presents t-SNE plots for the three remaining subsets in the GenImage dataset, namely the SDv1.5, Wukong and VQDM generators. As before, the top row shows embeddings extracted with the frozen ConvNeXt backbone (pre-LATTE) and the bottom row shows embeddings after LATTE fine-tuning.
Figure 8: Visualizations of t-SNE embeddings for real and fake images across the remaining three generators from GenImage. The first row presents embeddings before using LATTE (extracted using the original ConvNeXt), while the second row shows embeddings derived from LATTE. The reduced overlap in the second row indicates LATTE’s discriminative power and generalization capabilities.
Appendix D Latent Trajectory Spatial Analysis
To distinguish how diffusion-based reconstructions differ between real and generated images — and to motivate the modeling the latent trajectory — we analyze the spatial distribution of latent denoising corrections across timesteps.
Specifically, we compute the average per-pixel norm of latent differences between consecutive denoising steps - denoted as $\Delta z_{t}=|z_{t}-z_{t-1}|_{2}$ for a sequence of tracked timesteps ${t_{1},t_{2},\ldots,t_{K}}$. For each timestep interval $t_{k-1}\rightarrow t_{k}$, we aggregate $\Delta z_{t}$ across all spatial positions and across a batch of samples to obtain a mean spatial correction heatmap:
@@ Appendix D Latent Trajectory Spatial Analysis
By contrast, the ADM subset (9b) shows a markedly different trend. Here, the real vs. fake differences across all steps are considerably more subtle, and the resulting $\Delta z$ heatmaps for both classes appear visually similar in both magnitude and spatial pattern, with the exception of small brighter left and top margins. This behavior is consistent with our model’s relatively poor performance on ADM (74% compared to the 90.9% average) and suggests that the images in this subset are particularly difficult to distinguish - even in trajectory space.
Finally, SDv1.4 (9c) presents the most distinctive behavior. Unlike previous generators, the fake heatmaps exhibit a center-focused $\Delta z$ signature. This effect likely arises because we use Stable Diffusion [7] for both generating and reconstructing the images. The denoiser has learned to prioritize central content - where objects are typically located during prompt-guided generation - and thus applies larger, spatially focused corrections in the center of the image. Real images, by contrast, lack this learned structure and receive relatively uniform and lower-magnitude corrections across space.

### v2

@@ A.2 Effect of latent extraction configuration
Table 9: Accuracy(%) comparison between SDv1.5 and SDv2.1-based latent extraction. SDv2.1 improves accuracy by 3.9%.
U-Net Backbone Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
SDv1.5 81.7 99.7 99.5 77.7 90.6 98.0 78.1 71.5 87.1
@@
The results indicate that LATTE’s performance is sensitive to the latent extraction configuration, highlighting the importance of both the sampling method and the U-Net backbone. Switching from DDIM to DDPM yields a substantial improvement in average detection accuracy (+7.2%), with particularly large gains on datasets such as Midjourney and BigGAN. This suggests that the stochastic denoising dynamics captured by DDPM produce richer latent trajectories, enhancing the discriminative signal between real and generated images. Similarly, upgrading the U-Net backbone from SDv1.5 to SDv2.1 further improves average accuracy (+3.9%), reflecting the impact of more expressive latent representations on the model’s ability to capture subtle generative artifacts. While some datasets, such as ADM, show minimal changes, likely due to inherent detection difficulty or saturation effects, the overall trend confirms that both the scheduler and backbone play complementary roles: the scheduler shapes the temporal evolution of latents, whereas the backbone determines the quality of the underlying feature space. Despite these variations, LATTE maintains high and consistent performance across all configurations, demonstrating its robustness and reliability as a diffusion-generated image detector.
A.3 Influence of vision backbone fine-tuning
Our default setup fine-tunes the vision encoder. To quantify the added benefit of this choice, we compare against a variant where we freeze the backbone and train only the LATTE-specific components. Table 10 reports per-generator accuracy for both settings. We observe major improvements for both vision backbones when fine-tuned, with 15.5% accuracy gain for CLIP ViT-L/14 (Radford et al., 2021) and 9% for ConvNeXt (Liu et al., 2022). This likely stems from the fact that frozen backbones retain features that were never explicitly optimized for real vs. fake discrimination, leading to an embedding space that is misaligned with the objectives of generated image detection. Without adaptation, our model may struggle to effectively ground latent trajectories in meaningful visual semantics. Fine-tuning, by contrast, enables the backbone to specialize its representations for this task, enhancing the alignment between visual and latent features essential for robust detection.
Table 10: Accuracy (%) comparison for different vision backbones and fine-tuning vs. frozen settings on the GenImage dataset. Fine-tuned ConvNeXt yields the best performance.
@@ A.3 Influence of vision backbone fine-tuning
A.4 Influence of separate latent processing strategy
Table 11: Accuracy(%) comparison between separate vs. joint latent processing strategies. Processing timesteps separately yields the highest average accuracy, outperforming joint processing by 0.8%.
Joint 88.4 99.7 99.6 72.5 94.4 98.6 79.4 88.9 90.2
@@ A.4 Influence of separate latent processing strategy
A.5 Effect of positional encodings in CLS-pooling
We conduct an ablation to isolate the effect of the positional embeddings when using CLS-pooling. Specifically, we compare the full model (“CLS-pooling w/ pos. enc.”) to a variant that uses the same CLS-based self-attention but omits positional embeddings (“CLS-pooling w/o pos. enc.”), removing any explicit indication of timestep order. As shown in Table 12, providing sequence order information results in a significant improvement of 7.2% in average accuracy, confirming that timestep position is an important signal when aggregating latents jointly. Despite this gain, the CLS-based variant remains less effective than the default LATTE architecture, which aggregates the outputs via average pooling. Interestingly, the ”CLS-pooling w/ pos. enc.” variant demonstrates better performance on certain subsets - 9.2% increase on ADM and 7.9% on VQDM - suggesting that this CLS-based design, paired with sequence order cues, can be beneficial in specific contexts.
Table 12: Accuracy(%) comparison for CLS-pooling with and without explicit sequence order. Explicit positional embeddings improve accuracy by 7.2% over the implicit variant, but fall slightly short of the average pooling.
Sequence order Pos. enc. Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
CLS-pooling no 75.5 99.7 99.6 72.2 78.9 98.3 75.6 59.3 82.4
CLS-pooling yes 85.8 99.0 99.0 83.2 86.7 96.8 88.7 77.8 89.6
@@ A.5 Effect of positional encodings in CLS-pooling
Appendix B Latent Trajectory Spatial Analysis
To motivate the modeling of the latent trajectory and to distinguish how diffusion-based reconstructions differ between real and generated images, we analyze the spatial distribution of latent denoising corrections across timesteps.
Specifically, we compute the average per-pixel norm of latent differences between consecutive denoising steps - denoted as $\Delta z_{t}=|z_{t}-z_{t-1}|_{2}$ - for a sequence of tracked timesteps $t_{1},t_{2},\ldots,t_{K}$}. For each timestep interval $t_{k-1}\rightarrow t_{k}$, we aggregate $\Delta z_{t}$ across all spatial positions and across a batch of samples to obtain a mean spatial correction heatmap:
@@ Appendix B Latent Trajectory Spatial Analysis
By contrast, the ADM subset (7b) shows a markedly different trend. Here, the real vs. fake differences across all steps are considerably more subtle, and the resulting $\Delta z$ heatmaps for both classes appear visually similar in both magnitude and spatial pattern, with the exception of small brighter left and top margins. This behavior is consistent with our model’s relatively poor performance on ADM (74% compared to the 91% average) and suggests that the images in this subset are particularly difficult to distinguish - even in trajectory space.
Finally, SDv1.4 (7c) presents the most distinctive behavior. Unlike previous generators, the fake heatmaps exhibit a center-focused $\Delta z$ signature. This effect likely arises because we use Stable Diffusion (Rombach et al., 2022b) for both generating and reconstructing the images. The denoiser has learned to prioritize central content - where objects are typically located during prompt-guided generation - and thus applies larger, spatially focused corrections in the center of the image. Real images, by contrast, lack this learned structure and receive relatively uniform and lower-magnitude corrections across space.
Appendix C Complete accuracy & average precision on GenImage
Figure 8 presents LATTE’s performance on the GenImage dataset, reporting both accuracy and average precision across different training–testing generator combinations. The results show that LATTE maintains consistently high performance regardless of the generator used for training, highlighting its ability to generalize across diverse generative models.
Figure 8: Accuracy(%) (left) and average precision(%) (right) of LATTE across the GenImage dataset. The x-axis indicates the generator used to produce the training data, while each bar represents the model’s performance when tested on data from the different generators.
Appendix D Architectural Details of the CLS-pooling
We consider CLS-pooling as an alternative aggregation strategy (instead of average pooling), illustrated in Figure 9. After independently fusing each projected latent $\tilde{z}_{t}$ with visual features, through a stack of transformer decoder layers, a learnable token $z_{\text{CLS}}\in\mathbb{R}^{d}$ is prepended to the sequence of refined latent embeddings $\widetilde{\mathcal{T}}(x)=\{\tilde{z}_{t_{1}},\tilde{z}_{2},\dots,\tilde{z}_{t_{K}}\}$. Learnable positional embeddings are added to this sequence to inform the model of the order of timesteps. The sequence is then passed through a shared self-attention stack of transformer layers, allowing the CLS token $z_{\text{CLS}}$ to interact with the full latent trajectory and aggregate information across timesteps. The final CLS token output serves as the aggregated trajectory representation $\tilde{z}_{\text{agg}}$. The rest of the architecture remains the same as in Figure 2.
Figure 9: Overview of our proposed LATTE architecture with CLS pooling as an aggregation strategy (denoted in red). A learnable CLS token is prepended to the fused sequence and processed via a self-attention stack.
Appendix E Embedding Space Analysis
To complete our embedding space analysis from Section 4.4, Figure 10 presents t-SNE plots for the three remaining subsets in the GenImage dataset, namely the SDv1.5 (Rombach et al., 2022b), Wukong (Wukong, 2024), and VQDM (Gu et al., 2022) generators. As in Figure 4, the top row shows embeddings extracted with the frozen ConvNeXt backbone (Liu et al., 2022) (pre-LATTE) and the bottom row shows embeddings after LATTE fine-tuning. The much clearer separation in the second row illustrates LATTE’s discriminative power.
Figure 10: Visualizations of t-SNE embeddings for real and fake images across the remaining three generators from GenImage. The first row presents embeddings before using LATTE (extracted using the original ConvNeXt), while the second row shows embeddings derived from LATTE.